# Importance of accounting for student identities and intersectionality for creating equitable and inclusive physics learning environments

### v1

@@ IV Comparison to Previous Studies
We now situate this work by comparing it to previous qualitative studies related to experiences of women in physics such as Johnson’s [74] study at a small college and our previous study conducted in 2019 [23] at a large research university. This comparison focuses on analyzing all four domains of power and evaluating the relative experiences of female students in the physics learning environments. We believe the study discussed here at a small liberal arts college shows a physics learning environment which is significantly better than the one at the large research university [23] but not as equitable and inclusive as the environment at the small college in Johnson’s study [74], which can be taken as a gold standard that all physics departments should aspire to be like.
@@
First, we will discuss the interpersonal domain, which relates to individuals communicating with each other, or in our context, where students interact with peers and faculty members. In Johnson’s study [74], undergraduate women described their physics peers as “friendly,” “super nice,” and approachable. They perceived other physics students as partners in their learning so “why wouldn’t they help?” Students also got to know faculty members in their department. On the other hand, in Santana and Singh’s 2019 study [23], undergraduate women perceived their male peers to be condescending and to have many biases against women. In this current study, we saw that many women felt that their male peers were condescending and some male peers perceived these women to be incapable of doing physics without their help. Some of these women tried and managed to avoid their male peers. However, some of these women noted that they did not have extremely negative interactions with their male peers.
@@
In terms of the cultural domain, Johnson’s study showed an extremely positive physics culture, very atypical of many traditional physics department and classroom settings. Female students loved being around the physics building, thus showing that the physical space was welcoming and fostered an environment that emphasized working together. On the other hand, in our previous study [23], most women did not want to interact with male peers and the culture in spaces such as study groups or the undergraduate physics lounge, where men dominated, was very toxic. Both peers and instructors were reinforcing toxic and stereotypical physics cultures that included beliefs about physics only being for smart people, women not being able to do well in physics, etc. In this study, some of the women chose not to work with male peers. Specifically, Paulina felt that she could not work with anyone because of how they perceived her as a woman of color and due to the lack of strong commitment to work collaboratively even when she was placed in a group for the first time at the time the interviews were conducted. She also specifically felt that cliques were formed, especially if students already knew each other (e.g., from previous classes), so it was difficult for her to work effectively in groups. From Paulina’s perspective, we also saw how her identity as a woman of color impacted these interactions and she perceived her peers to negatively stereotype her, which further alienated her and made her anxious.
In Johnson’s study, she described the physics structural domain to be interactive and collaborative. For instance, faculty structured their classes to incorporate group work during lectures and faculty also guided students on how to work effectively in groups [74]. In our previous study, we noted that physics courses were taught traditionally, with few faculty incorporating active-learning strategies. In particular, faculty did not typically incorporate collaborative elements such as group work during their lectures. In this study, many of these women did describe working in groups, but did occasionally work alone. Thus, there is some evidence that some instructors had group work embedded in their lectures. However, many of them also called for more emphasis on peer collaboration by their instructors and effective norms for peer collaborations. For example, a few students suggested working in pairs as opposed to small groups in order to not be marginalized (e.g., Paulina) as well as to reduce cognitive load and get to know their peers well.
@@ V Summary, Conclusions and Future Directions
In this study, we interviewed seven undergraduate women in physics who attended a small liberal arts PWI. Through these interviews, we learned about their early experiences in physics during high school, how their family supported them in physics, e.g., by providing emotional and academic support, their interactions with their college peers and faculty, reflections about how certain intersections of their identity may have affected their experiences in physics, advice based upon what has helped them persist in physics, and suggestions for faculty that would improve their experiences.
@@
We note that these women’s family played a large role in supporting them, from an early age including during high school. They also provided a range of opportunities, which fostered interest in physics and science. However, beyond emotional support, most of these undergraduate women were unable to rely on family for academic support during college. They mainly credited some of their faculty as providing support for them. This might be the case as all of the women moved away from home to attend this college and their instructors and other physics faculty were closer in proximity to provide support especially in physics. Thus, based upon their experiences, instructors play a very important role in these women’s college experiences in physics once they are away from home.
@@
In particular, we see that their relationships with and perceptions of the faculty they regularly interacted with were generally overall positive. These women mentioned that women faculty were especially inspiring to them. Some of the faculty who served as academic advisors or instructors also encouraged a few of these women like Paulina or Ruby to major in physics. At least one of these faculty incorporated discussions about the experiences of marginalized people in physics. Despite these positive perceptions and experiences, some of these women noted that faculty often did not take actions or were not aware of equity and inclusion issues in their own classes. Thus, these undergraduate women had negative experiences with their peers which persisted without any intervention or with only superficial interventions by the instructor when such issues were explicitly brought out.
@@
Again, based upon domain of power and HELPIEE frameworks, we emphasize that physics instructors have a lot of power in their classrooms to empower students and so it is important that they use that power to address inequities in the classroom. For example, several of these women had experiences in which their male peers talked to them in a condescending way, took credit for their work, or in Paulina’s case, made her feel excluded from group work. Instructors appeared to be rarely aware of these experiences in their classroom or addressed such issues superficially when informed about them so that they persisted. Also, even if they communicated to their students that they could choose to switch partners, the women still felt like the onus was on them to bring up negative interactions with their peers to their instructor’s attention.
@@
We also saw that several women reflected on various aspects of their identities that affected their experiences negatively, which may be useful for faculty to consider. We encourage instructors to continue to contemplate upon and consider various intersections of students’ identity and how those may impact students in physics learning environments since students are not a monolith and have a wide range of backgrounds and experiences. Our findings could be useful useful for faculty professional develop workshop leaders in educating and informing instructors on how to better support diverse groups of students.
@@
In our interviews, We also asked these women about things that would help improve their experiences. Some of them provided advice based upon experiences that have personally helped them. However, we note that their advice may not be suitable for all students and some of them may put the onus on the marginalized students to change in order to improve their experiences, rather than focusing on the systemic issues that caused those negative experiences and the need for the physics culture to change to support them. Moreover, students who are marginalized in physics due to multiple marginalized identities, such as Paulina, may not benefit from simply being advised to change their mindset about their struggle unless the physics culture is improved to support them as some of the white women in this study suggested has helped them.
@@
These women also had many suggestions for instructors, many of which were about ways to change the structure of the classrooms. Others were more focused on the responsibility of the instructor. We encourage physics faculty to consider these suggestions as many of them do not require monetary resources, but require serious reflection and time commitment on their part. In other words, these suggestions are not impossible to implement, but may require planning and collaboration amongst faculty and other experts in positions of power. Faculty also do not need to be physics education researchers to carry out these changes; instead, they must be committed towards equity and inclusion.
@@
With regard to the domains of power and HELPIEE frameworks, the power in physics learning environments directly refers to the fact that physics faculty and other individuals can make the learning environments equitable and inclusive and take action and educate those who violate the equitable and inclusive norms. Enforcing rules, such as those described by a code of conduct, and even maintaining them through the repetitive use of discussions, training and disciplinary action can potentially be useful. However, at this small college where we conducted the interviews, from the perspective of these undergraduate women, the instructors were not aware of students not treating other students equally or if these issues were pointed out, they addressed these issues superficially in class as opposed to in a way that would make sustained difference. Thus, their actions in the disciplinary domain, if any, were relatively minor. The cultural domain in the domains of power framework is in reference to the physics culture, e.g., how members of the physics community think about physics, and how those beliefs are reinforced by other members. According to these undergraduate women, the physics culture at this college is much like a prototypical physics culture, e.g., where doing physics is associated with brilliance.
@@
As discussed in the preceding section, if we were to rank this physics department relative to the other two mentioned here [74, 23] in terms of the physics culture, it would be somewhere in the middle. However, at a small liberal arts institution, the culture may be easier to change because there are fewer faculty who have to be on the same page, whereas at a large university or a fragmented department, it would be more difficult to dismantle an established physics culture with, for example, 40+ physics faculty [23]. In such a large physics department with inequitable physics culture, there would need to be a “critical” number of faculty who are committed towards equity and inclusion in a physics department, who could work to change the physics culture [23]. With enough faculty support and resources, we believe that there can be a sustained and systemic effort in physics departments of any size and demographic distribution to create equitable and inclusive learning environments and support traditionally marginalized students (although some efforts could require more effort than others).
@@
In summary, through the experiences of undergraduate women presented here, we illustrated how they navigated and perceived the physics culture and engaged in doing physics. How power interacts in domains such as interpersonal, cultural, and disciplinary is especially important as they are entangled with each other. Because physics instructors and other administrators have large roles, influence, and power, according to the HELPIEE framework, they should take action, rather than expecting students to change or adapt to the physics culture. These actions should focus on traditionally marginalized students, e.g., they can focus on implementing interventions or workshops to dismantle stereotypes in the classroom. There is also need for systemic institutional changes. However, that would require more faculty and administrative support. In summary, we encourage physics faculty to take these findings and suggestions seriously and use their positions of power to implement positive changes and create a supportive environment for traditionally marginalized students.
Acknowledgements.

### v2

@@ V.3 Comparing the three studies
This comparison focuses on analyzing all four Domains of Power and evaluating the relative experiences of women students in the physics learning environments. We believe the study discussed here in a medium-size physics department at a small liberal arts college shows a physics learning environment which is significantly better than the one at the large research university [23] but not as equitable and inclusive as the environment in the small physics department in Johnson’s study [70], which can be taken as a gold standard that all physics departments should aspire to be like. Thus, these two previous studies are useful for making comparisons.
@@
First, we will discuss the interpersonal domain, which relates to individuals communicating with each other, or in our context, where students interact with peers and faculty members. In Johnson’s study [70], undergraduate women described their physics peers as “friendly,” “super nice,” and approachable. They perceived other physics students as partners in their learning so “why wouldn’t they help?” Students also got to know faculty members in their department. On the other hand, in Santana and Singh’s 2019 study [23], undergraduate women perceived their male peers to be condescending and to have many biases against women. In this current study, we saw that many women felt that their male peers were condescending and some male peers perceived the women to be incapable of doing physics without their help. Some of the women tried and managed to avoid their male peers. However, some of the women noted that they did not have extremely negative interactions with their male peers.
@@
In terms of the cultural domain, Johnson’s study showed an extremely positive physics culture, very atypical of many traditional physics department and classroom settings. Women students loved being around the physics building, thus showing that the physical space was welcoming and fostered an environment that emphasized working together. On the other hand, in our previous study [23], most women did not want to interact with male peers and the culture in spaces such as study groups or the undergraduate physics lounge, where men dominated, was very toxic. Both peers and instructors were reinforcing toxic and stereotypical physics cultures that included beliefs about physics only being for smart people, women not being able to do well in physics, etc. In this study, some of the women chose not to work with male peers. Specifically, Paulina felt that she could not work with anyone because of how they perceived her as a woman of color and due to the lack of strong commitment to work collaboratively even when she was placed in a group for the first time at the time the interviews were conducted. She also specifically felt that cliques were formed, especially if students already knew each other (e.g., from previous classes), so it was difficult for her to work effectively in groups. From Paulina’s perspective, we also saw how her identity as a woman of color impacted these interactions and she perceived her peers to negatively stereotype her, which further alienated her and made her anxious.
In Johnson’s study, she described the physics structural domain to be interactive and collaborative. For instance, faculty structured their classes to incorporate group work during lectures and faculty also guided students on how to work effectively in groups [70]. In our previous study, we noted that physics courses were taught traditionally, with few faculty incorporating active-learning strategies. In particular, faculty did not typically incorporate collaborative elements such as group work during their lectures. In this study, many of the women did describe working in groups, but did occasionally work alone. Thus, there is some evidence that some instructors had group work embedded in their lectures. However, many of them also called for more emphasis on peer collaboration by their instructors and effective norms for peer collaborations. For example, a few students suggested working in pairs as opposed to small groups in order to not be marginalized (e.g., Paulina) as well as to reduce cognitive load and get to know their peers well.
@@ VI Conclusions
In this study, we interviewed seven undergraduate women majoring in physics in a medium-size physics department who attended a small liberal arts PWI in the US. Through these interviews, we learned about their early experiences in physics during high school, how their family supported them in physics, e.g., by providing emotional and academic support, their interactions with their college peers and faculty, reflections about how certain intersections of their identity may have affected their experiences in physics, advice based upon what has helped them persist in physics, and suggestions for faculty that would improve their experiences.
@@
We note that the women’s family played a large role in supporting them, from an early age including during high school. They also provided a range of opportunities, which fostered interest in physics and science. However, beyond emotional support, most of these undergraduate women were unable to rely on family for academic support during college. They mainly credited some of their faculty as providing support for them. This might be the case as all of the women moved away from home to attend this college and their instructors and other physics faculty were closer in proximity to provide support especially in physics. Thus, based upon their experiences, instructors play a very important role in the women’s college experiences in physics once they are away from home.
@@
In particular, we see that their relationships with and perceptions of the faculty they regularly interacted with were generally overall positive. The women mentioned that women faculty were especially inspiring to them. Some of the faculty who served as academic advisors or instructors also encouraged a few of the women like Paulina or Ruby to major in physics. At least one of these faculty incorporated discussions about the experiences of marginalized people in physics. Despite these positive perceptions and experiences, some of the women noted that faculty often did not take actions or were not aware of equity and inclusion issues in their own classes. Thus, these undergraduate women had negative experiences with their peers which persisted without any intervention or with only superficial interventions by the instructor when such issues were explicitly brought out.
@@
Again, based upon Domains of Power and HELPIEE frameworks, we emphasize that physics instructors have a lot of power in their classrooms to empower students and so it is important that they use that power to address inequities in the classroom. For example, several of the women had experiences in which their male peers talked to them in a condescending way, took credit for their work, or in Paulina’s case, made her feel excluded from group work. Instructors appeared to be rarely aware of these experiences in their classroom or addressed such issues superficially when informed about them so that they persisted. Also, even if they communicated to their students that they could choose to switch partners, the women still felt like the onus was on them to bring up negative interactions with their peers to their instructor’s attention.
@@
We also saw that several women reflected on various aspects of their identities that affected their experiences negatively, which may be useful for faculty to consider. We encourage instructors to continue to contemplate upon and consider various intersections of students’ identity and how those may impact students in physics learning environments since students are not a monolith and have a wide range of backgrounds and experiences. Our findings could be useful for faculty professional development workshop leaders in educating and informing instructors on how to better support diverse groups of students.
@@
In our interviews, we also asked the women about things that would help improve the physics learning environments. Some of them provided advice based upon experiences that have personally helped them. However, we note that their advice may not be suitable for all students and some of them may put the onus on the marginalized students to change in order to improve their experiences, rather than focusing on the systemic issues that caused those negative experiences and the need for the physics culture to change to support them. Moreover, students who are marginalized in physics due to multiple marginalized identities, such as Paulina, may not benefit from simply being advised to change their mindset about their struggle unless the physics culture is improved to support them as some of the white women in this study suggested has helped them.
@@
The women also had many suggestions for instructors, many of which were about ways to change the structure of the classrooms. Others were more focused on the responsibility of the instructor. We encourage physics faculty to consider these suggestions as many of them do not require monetary resources, but require serious reflection and time commitment on their part. In other words, these suggestions are not impossible to implement, but may require planning and collaboration amongst faculty and other experts in positions of power. Faculty also do not need to be physics education researchers to carry out these changes; instead, they must be committed towards equity and inclusion. We also note that because instructors have power in several domains, they are responsible for setting the tone for how students are treated and establish classroom norms. In addition, for many physics programs, men make up the majority of both physics majors and physics faculty. Hence, they have both power and influence to make change.This change is necessary and being neutral signals both a privilege and polite oppression, which is a phrase coined by Ferretti [81].
@@
With regard to the Domains of Power and HELPIEE frameworks, the power in physics learning environments directly refers to the fact that physics faculty and other individuals can make the learning environments equitable and inclusive and take action and educate those who violate the equitable and inclusive norms. Enforcing rules, such as those described by a code of conduct, and even maintaining them through the repetitive use of discussions, training and disciplinary action can potentially be useful. However, at this small college where we conducted the interviews, from the perspective of these undergraduate women, the instructors were not aware of students not treating other students equally or if these issues were pointed out, they addressed these issues superficially in class as opposed to in a way that would make sustained difference. Thus, their actions in the disciplinary domain, if any, were relatively minor. The cultural domain in the Domains of Power framework pertains, e.g., to how members of the physics community think about physics, and how those beliefs are reinforced by other members. According to these undergraduate women, the physics culture at this college is much like a prototypical physics culture, e.g., where doing physics is associated with brilliance.
@@
As discussed in the preceding section, if we were to rank this physics department relative to the other two mentioned here [70, 23] in terms of the physics culture, it would be somewhere in the middle. However, at a small liberal arts institution with medium-size or small physics department, the culture may be easier to change because there are fewer faculty who have to be on the same page, whereas at a large research university or a fragmented department, it would be more difficult to dismantle an established physics culture with, for example, 40+ physics faculty [23]. In such a large physics department with inequitable physics culture, there would need to be a “critical” number of faculty who are committed towards equity and inclusion in a physics department, who could work to change the physics culture [23]. With enough faculty support and resources, we believe that there can be a sustained and systemic effort in physics departments of any size and demographic distribution to create equitable and inclusive learning environments and support traditionally marginalized students, although some initiatives could require more effort than others.
@@
In summary, through the experiences of undergraduate women presented here, we illustrated how they navigated and perceived the physics culture and engaged in doing physics. How power interacts in domains such as interpersonal, cultural, and disciplinary is especially important as they are entangled with each other. Because physics instructors and other administrators have large roles, influence, and power, according to the HELPIEE framework, they should take action, rather than expecting students to change or adapt to the physics culture. These actions should focus on traditionally marginalized students, e.g., they can focus on implementing interventions or workshops to dismantle stereotypes in the classroom. There is also need for systemic institutional changes that would require more faculty and administrative support. In summary, physics faculty should take these findings and suggestions seriously and use their positions of power to implement positive changes and create a supportive environment for traditionally marginalized students.
VII Limitations and Future Directions
One limitation of the study is that participants, although they constituted one third of the women physics majors at the time, were self-selected, i.e., they volunteered. This may have limited other types of experiences from other women in the physics program at this institution. Future work can investigate women’s experiences at other physics departments of different sizes at various types of institutions including minority serving institutions. This can also ensure that their perspectives are represented in the broader physics education research community. It would also be valuable to carry out similar investigations in other countries with undergraduate women majoring in physics.
We also note that the protocol questions were not explicitly shaped around the Domains of Power framework. However, many of the questions aligned with both Standpoint Theory and the Domains of Power frameworks. For example, we asked women questions regarding their interactions with peers and faculty, class structure, and physics culture. In answering the protocol and follow-up questions, the women reflected on the instructors’ actions which can be mapped onto different Domains of Power. We also note that Johnson used Domains of Power framework to analyze data in her 2020 study [70] and asked questions similar to our questions instead of asking explicit questions shaped around the Domains of Power framework.
Acknowledgements
@@
VIII Appendix
VIII.1 Protocol Questions
Table 2: Example of the relevant protocol questions used to probe undergraduate women’s experiences.
Example Protocol Questions
Can you tell me about your experiences with science, math, and physics in high school? How does college science compare to learning in high school?
What contributed to your decision to major in physics?
Did you work with other people on homework, or studying for tests, in your first year in physics? What about other classes?
Do you typically work with people of the same gender as you, a different gender, or a mixture? Which do you prefer?
Can you describe an occasion where you struggled in the physics program? How do you react when you face challenges like this?
Do you feel that your contributions are valued by your peers and instructors?
Do you feel respected as a physics person by your peers and instructor?
Have you had any mentors or people who inspired you to do physics, in school, at college, or in your outside life?
Do you feel supported by your instructors in your physics courses?
Do you think your experiences in physics have been different because of your gender? If so, how?
What advice would you give to young women who are considering majoring in physics?
What could the physics department do to make the environment less chilly, or to remove barriers, or to make the playing field level for men and women?
VIII.2 General codes
Figure 3: Examples of general codes from participant data. Each color refers to a separate theme. The key lists abbreviations in some cases. Some of these themes were not discussed in this paper.
VIII.3 Definitions and examples for general codes.
Table 3: Section numbers for analytic themes (AT), subthemes, and corresponding definitions and examples for general codes. Analytic themes are located in the main document (section IV: Findings and Discussion).
AT Subtheme General Code Definition Example from Data
A Family’s role in providing support Family providing academic support Ways in which family provided insight about science or help with homework “whenever I had issues with like homework and wanted help, I would always…go to her and she would help me out…”
A Family’s role in providing support Family providing exposure to science Ways in which family members, whose jobs were related to STEM, would discuss science topics, bring them to their place of work, or create a positive experience related to STEM “…my grandpa was very into science and at a young age, he would always do these math puzzles with me and I just found [it] really interesting…”
A Positive high school experiences Positive interactions with teachers Interactions students had with their teachers that were positive. For example, teachers were supportive, kind, helpful, etc. “I always felt they treated me like I was smart and capable and deserved to be there…”
A Positive high school experiences Positive interactions with peers Interactions students had with their peers that were positive. For example, peers were inspirational, driven, etc. “…it was nice being in a class where I had friends there and then they’re all very driven.”
A Negative high school experiences Negative interactions with teachers Interactions students had with their teachers that were negative. For example, teachers showed biases towards boys’ opinions, did not focus on learning, or were unprofessional. “…even if my hand was up or if another girl’s hand [was]…[my teacher] would always call on the male students before the female students”
A Negative high school experiences Negative interactions with peers Interactions students had with their peers that were negative. For example, peers were focused on grades as opposed to learning. “…I felt like everyone around me was thinking a lot about GPA and not really caring about the learning so I got a little discouraged..”
B Positive interactions & perceptions Positive experiences with college peers Interactions and opinions students had with their peers where they felt supported, safe, comfortable, working with them. “…people I really talk to are other girls and that has been the most comfortable for me”
B Negative interactions & perceptions Negative experiences with college peers Interactions and opinions students had with their peers where men dominated conversations, acted condescending, etc. “…if me and my partner had different answers, it kind of felt like he’d automatically assume that he was right, which wasn’t always the case…”
B Negative interactions & perceptions Negative opinions of college peers Negative opinions of male peers. Women believed men thought they were more capable or smarter than them. “…I think he thought of women in a sense, that they were less capable and more vulnerable than men, and so he wanted to be this savior…”
B Paulina as an island among her peers Being negatively perceived by peers Instances where Paulina reflects on how her intersectional identity impacts her interactions with peers. “there is more of a willingness to either poke fun or underestimate me because of the intersectionality of my gender and race.”
B Paulina as an island among her peers Being ignored by peers Instances where peers do not listen to or ignore Paulina. Other students choose not to engage with her. “…so a lot of the times I’ll bring [up] points and it’ll just be ignored and that’s something that I’ve had to deal with a lot.”
B Paulina as an island among her peers Working alone Caused by being ignored by others. These include instances, e.g., where Paulina feels like working alone is the only option . “…no one is reaching out to each other to talk and work things out together so by default I just feel like I have to end up doing things on my own.”
C Positive interactions & perceptions Positive interactions with instructors Positive interactions students had with instructors. This includes things instructors did that students thought were helpful. “[my instructors] have really made attempts to make their classroom environments better…”
C Positive interactions & perceptions Positive opinions about instructors Instances when students have positive opinions about instructors. For example, they thought their instructors were caring, helpful, supportive. “…we’re about to do a project and she’s trying to make sure not to put the women in our class as the only woman in a group…and I definitely am grateful for that…”
C Negative interactions & perceptions Negative interactions with instructors Negative interactions students had with instructors, including being singled out. “…the instructor has kind of singled me out in the classroom and asked me questions continuously trying to…make sure that I’m keeping up with the class…”
C Negative interactions & perceptions Negative opinions about instructors Instances when students have negative opinions about instructors. For example, they thought instructors were not aware of what was going on in the classroom. “…a lot of professors don’t go out of their way to look at the problems and…will assume [things are fine] or [don’t] want to have to deal with anything like that.”
D Reflection about identities  [^3] This is also the Analytic Theme since there is only one subtheme. Reflection on gender Moments where women reflected on how their gender identity impacted their experiences in physics. “When there’s less women, I feel some of that pressure…to like be smart, if you feel like you’re representing woman…”
D Reflection about identities Reflection on ethnicity/race Moments where women reflected on how their ethnic/racial identity impacted their experiences in physics. “there’s also been this expectation that I am just not up to par with them, be- cause [of] my racial background…”
D Reflection about identities Reflection on personality Moments where women reflected on how their personality impacted their experiences in physics. “…I tend to be more introverted…outside of the classroom. I think personality does create some divides…”
D Reflection about identities Reflection on education Moments where women reflected on how their prior education impacted their experiences in physics. “I think I didn’t so much feel like isolated or alienated because of that…I think again, I’ve been pretty cushioned in physics.”
E People providing emotional support Support from Family Ways in which family members provided emotional support or encouragement during their undergraduate degree. “They’d all told me, ‘You gotta go with your gut… because you know what you can do and you can do more than what you think you can do,’ and that really stuck with me.”
E People providing emotional support Support from Faculty Members Ways in which faculty members provided emotional support or encouragement during their physics trajectory. “…having someone that took the time to have a conversation with me and help me figure things out was really meaningful to me.”
E People providing emotional support Faculty as Role Models Ways in which women looked up to their instructors or were inspired by them. “He’s one of the few other Latinx professors that I’ve seen on campus… it’s become very meaningful to see someone from a similar background as me kind of succeed.”
F Personal Advice Involvement in the Department Ways in which women became involved in the physics department through leadership roles. These roles are ways that have provided a better experience for the women. “…I liked that once I was involved, I could then kind of make the events that I thought would be helpful so I’m glad I did that.”
F Personal Advice Persisting with a Growth Mindset Ways in which women adopted a growth mindset and how it has helped them persist in the physics major. “…what has helped me the most is learning that it’s okay to not understand everything… it’s about just practicing and trying and also just learning as much as you can.”
F Advice for Instructors Avoid forming opinions based on students’ backgrounds Suggestion provided by women for faculty members not to make assumptions about students based on their identity or background. “…a small school like [my college], where you might have professors for multiple years, you might come back to that professor that you’ve had your first semester and I think that’s important.. for professors to know.”
F Advice for Instructors Address and dismantle biases against women Suggestion provided by women for faculty members to communicate and address stereotypes against women in science “…making it a very, very clear fact that…men are no better at physics than women.”
F Advice for Instructors Address lack of representation of women in physics. Suggestion provided by women for faculty members to address the historical lack of representation of women in physics. “I think it’d be great to address the fact that there are not many historical women in physics, but that does not mean that there cannot be in the future…”
F Advice for Instructors Provide examples of contribution of women in physics Suggestion provided by women for faculty members to provide examples of contributions of women and need for more support for women in physics. “…seeing more women just in their discoveries is something that probably can be helpful.”
F Advice for Instructors Provide reassurance during struggle Suggestion provided by women for faculty members to recognize and reassure students regarding their performance and struggle. “…making it more clear that it’s okay to not always get the right answer…and that doesn’t mean you’re stupid…”
F Advice for Instructors Improve office hours Suggestion provided by women for faculty members to improve the structure and culture of office hours. “…get rid of this notion that students are supposed to come in already knowing everything…that they’re asking about.”
F Advice for Instructors Encourage peer collaboration Suggestion provided by women for faculty members to encourage students to work together and form community. “…I’d love to be able to talk with other students in a classroom as well as outside the classroom and learn from them and I don’t think there’s enough of that.”
F Advice for Instructors Restructure in-class groups Suggestion provided by women for faculty members to change how groups are formed. “…having rotations where people aren’t always stuck with the same group but also have enough time working with other people to break down these barriers..”
F Advice for Instructors Reinforce disciplinary actions Suggestion for instructors to discipline students when they are not working equitably. “if they’re not aware of [issues in the classroom], there’s no way for them to personally intervene….”
F Advice for Instructors Check-in with Students Suggestion for instructors to check on the well-being of students. “asking students how they feel in the space and [whether] they feel comfortable, they feel their questions are being answered…what [are] their interactions with students…”
F Advice for Instructors Engage in conversations with students about inequities Suggestion for instructors to have conversations with students about inequities in physics. “…having conversations where…students and professors would engage with the fact that, there is a lack of equity in the classroom for students of color…”